# Ultrafast Band Structure Dynamics in Bulk 1T-VSe

### v1

@@ III.2 Pump-probe spectroscopy
However, in our case of incresed intensity below the Fermi level, this would mean that the defect levels are not completely occupied under equilibrium conditions.
@@
An additional explanation of the metastable signal can be related to the light-induced suppression of electronic correlation in 1T-VSe, resulting in the recovery of additional density of states in the energy region of the vanadium 3d bands. Indeed, Biswas et al.[9] reported that significant electronic correlation is present in VSe. In particular, the fingerprint of this effect was found in the large broadening of the vanadium 3d states below the Fermi level, which could only be captured by dynamical mean field theory calculations with a large Hubbard interaction strength of the order eV. Similarly to what reported in [31], when energy is provided to the system via photoexcitation, the strength of electronic correlations can be markedly reduced, hence new electronic states in the energy region of the vanadium 3d bands, that were suppressed by localization, can emerge, similarly to the collapse of a correlation gap in photoexcited Mott insulators.
@@
In this study we used polarization-dependent ARPES to disentangle the contributions of different vanadium- and selenium-derived bands to the valence band structure close to the and A points of the 1T-VSe hexagonal BZ. Our results show good agreement with theoretical band structure calculations, allowing for one-to-one assignments of the probed bands. Furthermore, we used our extracted knowledge about the position and dispersion of different bands for interpreting the changes in the electronic density of states after optical pumping. We detect the formation of a metastable configuration of the density of states reached after optical pumping, leading to additional electron density of states spread over 0.5 eV below the Fermi-level, which has not been observed so far and that asks for further investigation. However, the possibility to alter the electronic bands by ultrashort light pulses in a way not attainable under equilibrium paves the way to the design of new metastable states of matter with tailored electronic properties and novel functionalities, connected to the light-induced enhanced metallicity of the system, under the form of an enhanced density of states extending from the valence-band region up to the Fermi level.
@@ Acknowledgements.
W.B. acknowledges financial support from the German academic exchange service (DAAD) for financing a research stay at Academia Sinica in Taiwan for establishing the collaboration.References
- Manzeli et al. [2017] S. Manzeli, D. Ovchinnikov, D. Pasquier, O. V. Yazyev, and A. Kis, Nature Reviews Materials 2, 17033 (2017).
@@
- Morrison et al. [2014] V. R. Morrison, R. P. Chatelain, K. L. Tiwari, A. Hendaoui, A. Bruhács, M. Chaker, and B. J. Siwick, Science 346, 445 (2014).
@@
- Muraoka et al. [2018] Y. Muraoka, H. Nagao, Y. Yao, T. Wakita, K. Terashima, T. Yokoya, H. Kumigashira, and M. Oshima, Scientific Reports 8, 17906 (2018).
@@
- Kogar et al. [2020] A. Kogar, A. Zong, P. E. Dolgirev, X. Shen, J. Straquadine, Y.-Q. Bie, X. Wang, T. Rohwer, I.-C. Tung, Y. Yang, R. Li, J. Yang, S. Weathersby, S. Park, M. E. Kozina, E. J. Sie, H. Wen, P. Jarillo-Herrero, I. R. Fisher, X. Wang, and N. Gedik, Nature Physics 16, 159 (2020).
@@
- Hedayat et al. [2019] H. Hedayat, C. J. Sayers, D. Bugini, C. Dallera, D. Wolverson, T. Batten, S. Karbassi, S. Friedemann, G. Cerullo, J. van Wezel, S. R. Clark, E. Carpene, and E. Da Como, Phys. Rev. Res. 1, 023029 (2019).
@@
- Duan et al. [2023] S. Duan, W. Xia, C. Huang, S. Wang, L. Gu, H. Liu, D. Xiang, D. Qian, Y. Guo, and W. Zhang, Phys. Rev. Lett. 130, 226501 (2023).
@@
- Shi et al. [2019] X. Shi, W. You, Y. Zhang, Z. Tao, P. M. Oppeneer, X. Wu, R. Thomale, K. Rossnagel, M. Bauer, H. Kapteyn, and M. Murnane, Science Advances 5, eaav4449 (2019).
@@
- Majchrzak et al. [2021] P. Majchrzak, S. Pakdel, D. Biswas, A. J. H. Jones, K. Volckaert, I. Marković, F. Andreatta, R. Sankar, C. Jozwiak, E. Rotenberg, A. Bostwick, C. E. Sanders, Y. Zhang, G. Karras, R. T. Chapman, A. Wyatt, E. Springate, J. A. Miwa, P. Hofmann, P. D. C. King, N. Lanatà, Y. J. Chang, and S. Ulstrup, Phys. Rev. B 103, L241108 (2021).
@@
- Biswas et al. [2021] D. Biswas, A. J. H. Jones, P. Majchrzak, B. K. Choi, T.-H. Lee, K. Volckaert, J. Feng, I. Marković, F. Andreatta, C.-J. Kang, H. J. Kim, I. H. Lee, C. Jozwiak, E. Rotenberg, A. Bostwick, C. E. Sanders, Y. Zhang, G. Karras, R. T. Chapman, A. S. Wyatt, E. Springate, J. A. Miwa, P. Hofmann, P. D. C. King, Y. J. Chang, N. Lanatà, and S. Ulstrup, Nano Letters 21, 1968 (2021).
- Umemoto et al. [2019] Y. Umemoto, K. Sugawara, Y. Nakata, T. Takahashi, and T. Sato, Nano Research 12, 165 (2019).
@@
- Feng et al. [2018] J. Feng, D. Biswas, A. Rajan, M. D. Watson, F. Mazzola, O. J. Clark, K. Underwood, I. Marković, M. McLaren, A. Hunter, D. M. Burn, L. B. Duffy, S. Barua, G. Balakrishnan, F. Bertran, P. Le Fèvre, T. K. Kim, G. van der Laan, T. Hesjedal, P. Wahl, and P. D. C. King, Nano Letters 18, 4493 (2018).
@@
- Duvjir et al. [2018] G. Duvjir, B. K. Choi, I. Jang, S. Ulstrup, S. Kang, T. Thi Ly, S. Kim, Y. H. Choi, C. Jozwiak, A. Bostwick, E. Rotenberg, J.-G. Park, R. Sankar, K.-S. Kim, J. Kim, and Y. J. Chang, Nano Letters 18, 5432 (2018).
@@
- Diego et al. [2021] J. Diego, A. H. Said, S. K. Mahatha, R. Bianco, L. Monacelli, M. Calandra, F. Mauri, K. Rossnagel, I. Errea, and S. Blanco-Canosa, Nature Communications 12, 598 (2021).
- Terashima et al. [2003] K. Terashima, T. Sato, H. Komatsu, T. Takahashi, N. Maeda, and K. Hayashi, Phys. Rev. B 68, 155108 (2003).
@@
- Strocov et al. [2012] V. N. Strocov, M. Shi, M. Kobayashi, C. Monney, X. Wang, J. Krempasky, T. Schmitt, L. Patthey, H. Berger, and P. Blaha, Phys. Rev. Lett. 109, 086401 (2012).
@@
- Fanfarillo et al. [2021] L. Fanfarillo, D. Kopić, A. Sterzi, G. Manzoni, A. Crepaldi, D. T. Payne, W. Bronsch, V. Tsurkan, D. Croitori, J. Deisenhofer, F. Parmigiani, M. Capone, and F. Cilento, Journal of Electron Spectroscopy and Related Phenomena 250, 147090 (2021).
@@
- Peli et al. [2020] S. Peli, D. Puntel, D. Kopic, B. Sockol, F. Parmigiani, and F. Cilento, Journal of Electron Spectroscopy and Related Phenomena 243, 146978 (2020).
- Kresse and Hafner [1993] G. Kresse and J. Hafner, Phys. Rev. B 48, 13115 (1993).
- Kresse and Furthmüller [1996a] G. Kresse and J. Furthmüller, Computational Materials Science 6, 15 (1996a).
- Kresse and Furthmüller [1996b] G. Kresse and J. Furthmüller, Phys. Rev. B 54, 11169 (1996b).
- Blöchl [1994] P. E. Blöchl, Phys. Rev. B 50, 17953 (1994).
- Kresse and Joubert [1999] G. Kresse and D. Joubert, Phys. Rev. B 59, 1758 (1999).
- Perdew et al. [1996] J. P. Perdew, K. Burke, and M. Ernzerhof, Phys. Rev. Lett. 77, 3865 (1996).
@@
- Jolie et al. [2019] W. Jolie, T. Knispel, N. Ehlen, K. Nikonov, C. Busse, A. Grüneis, and T. Michely, Phys. Rev. B 99, 115417 (2019).
@@
- Falke et al. [2021] Y. Falke, N. Ehlen, G. Marini, A. V. Fedorov, V. Y. Voroshnin, B. V. Senkovskiy, K. Nikonov, M. Hoesch, T. K. Kim, L. Petaccia, G. Di Santo, T. Szkopek, G. Profeta, and A. Grüneis, Phys. Rev. B 104, 235137 (2021).
- Henke et al. [2020] J. Henke, F. Flicker, J. Laverock, and J. van Wezel, SciPost Phys. 9, 56 (2020).
@@
- Tamura et al. [1985] E. Tamura, R. Feder, J. Krewer, R. Kirby, E. Kisker, E. Garwin, and F. King, Solid State Communications 55, 543 (1985).
- Claessen et al. [1990] R. Claessen, I. Schafer, and M. Skibowski, Journal of Physics: Condensed Matter 2, 10045 (1990).
- Yilmaz and Vescovo [2022] T. Yilmaz and E. Vescovo, Phys. Rev. B 105, L041101 (2022).
- Wang et al. [2015] H. Wang, C. Zhang, and F. Rana, Nano Letters 15, 339 (2015).
@@
- Cilento et al. [2018] F. Cilento, G. Manzoni, A. Sterzi, S. Peli, A. Ronchi, A. Crepaldi, F. Boschini, C. Cacho, R. Chapman, E. Springate, H. Eisaki, M. Greven, M. Berciu, A. F. Kemper, A. Damascelli, M. Capone, C. Giannetti, and F. Parmigiani, Science Advances 4, eaar1998 (2018).

### v2

@@ III.2 Pump-probe spectroscopy
In our case of increased intensity below the Fermi level, this would mean that the defect levels are dragged below the Fermi level upon population.
@@
An additional explanation of the metastable signal can be related to the light-induced suppression of electronic correlation in 1T-VSe2, resulting in the recovery of additional density of states in the energy region of the vanadium 3d bands. Indeed, Biswas et al.Biswas et al. (2021) reported that significant electronic correlation is present in VSe2. In particular, the fingerprint of this effect was found in the large broadening of the vanadium 3d states below the Fermi level, which could only be captured by dynamical mean field theory calculations with a large Hubbard interaction strength of the order eV. Similarly to what reported in Cilento et al. (2018), when energy is provided to the system via photoexcitation, the strength of electronic correlations can be markedly reduced, hence new electronic states in the energy region of the vanadium 3d bands, that were suppressed by localization, can emerge, similarly to the collapse of a correlation gap in photoexcited Mott insulators.
@@
In this study we used polarization-dependent ARPES to disentangle the contributions of different vanadium- and selenium-derived bands to the valence band structure close to the and A points of the 1T-VSe2 hexagonal BZ. Our results show good agreement with theoretical band structure calculations, allowing for one-to-one assignments of the probed bands. Furthermore, we used our extracted knowledge about the position and dispersion of different bands to interpret the changes in the electronic density of states out-of-equilibrium. We detect the formation of a metastable configuration of the density of states reached after optical pumping, leading to additional electron density of states spread over 0.5 eV below the Fermi level, which has not been observed so far. This finding asks for further investigation. Nonetheless, the possibility to alter the electronic bands by ultrashort light pulses in a way not attainable under equilibrium paves the way to the design of new metastable states of matter with tailored electronic properties and novel functionalities, connected to the light-induced enhanced metallicity of the system, under the form of an enhanced density of states extending from the valence-band region up to the Fermi level.
@@ Acknowledgements.
W.B. acknowledges financial support from the German Academic Exchange Service (DAAD) for financing a research stay at the Institute of Atomic and Molecular Sciences at Academia Sinica in Taipei (Taiwan) for establishing collaborations.References
- Manzeli et al. (2017) S. Manzeli, D. Ovchinnikov, D. Pasquier, O. V. Yazyev, and A. Kis, Nature Reviews Materials 2, 17033 (2017).
@@
- Morrison et al. (2014) V. R. Morrison, R. P. Chatelain, K. L. Tiwari, A. Hendaoui, A. Bruhács, M. Chaker, and B. J. Siwick, Science 346, 445 (2014).
@@
- Muraoka et al. (2018) Y. Muraoka, H. Nagao, Y. Yao, T. Wakita, K. Terashima, T. Yokoya, H. Kumigashira, and M. Oshima, Scientific Reports 8, 17906 (2018).
@@
- Kogar et al. (2020) A. Kogar, A. Zong, P. E. Dolgirev, X. Shen, J. Straquadine, Y.-Q. Bie, X. Wang, T. Rohwer, I.-C. Tung, Y. Yang, R. Li, J. Yang, S. Weathersby, S. Park, M. E. Kozina, E. J. Sie, H. Wen, P. Jarillo-Herrero, I. R. Fisher, X. Wang, and N. Gedik, Nature Physics 16, 159 (2020).
@@
- Hedayat et al. (2019) H. Hedayat, C. J. Sayers, D. Bugini, C. Dallera, D. Wolverson, T. Batten, S. Karbassi, S. Friedemann, G. Cerullo, J. van Wezel, S. R. Clark, E. Carpene, and E. Da Como, Phys. Rev. Res. 1, 023029 (2019).
@@
- Duan et al. (2023) S. Duan, W. Xia, C. Huang, S. Wang, L. Gu, H. Liu, D. Xiang, D. Qian, Y. Guo, and W. Zhang, Phys. Rev. Lett. 130, 226501 (2023).
@@
- Shi et al. (2019) X. Shi, W. You, Y. Zhang, Z. Tao, P. M. Oppeneer, X. Wu, R. Thomale, K. Rossnagel, M. Bauer, H. Kapteyn, and M. Murnane, Science Advances 5, eaav4449 (2019).
@@
- Majchrzak et al. (2021) P. Majchrzak, S. Pakdel, D. Biswas, A. J. H. Jones, K. Volckaert, I. Marković, F. Andreatta, R. Sankar, C. Jozwiak, E. Rotenberg, A. Bostwick, C. E. Sanders, Y. Zhang, G. Karras, R. T. Chapman, A. Wyatt, E. Springate, J. A. Miwa, P. Hofmann, P. D. C. King, N. Lanatà, Y. J. Chang, and S. Ulstrup, Phys. Rev. B 103, L241108 (2021).
@@
- Biswas et al. (2021) D. Biswas, A. J. H. Jones, P. Majchrzak, B. K. Choi, T.-H. Lee, K. Volckaert, J. Feng, I. Marković, F. Andreatta, C.-J. Kang, H. J. Kim, I. H. Lee, C. Jozwiak, E. Rotenberg, A. Bostwick, C. E. Sanders, Y. Zhang, G. Karras, R. T. Chapman, A. S. Wyatt, E. Springate, J. A. Miwa, P. Hofmann, P. D. C. King, Y. J. Chang, N. Lanatà, and S. Ulstrup, Nano Letters 21, 1968 (2021).
- Umemoto et al. (2019) Y. Umemoto, K. Sugawara, Y. Nakata, T. Takahashi, and T. Sato, Nano Research 12, 165 (2019).
@@
- Feng et al. (2018) J. Feng, D. Biswas, A. Rajan, M. D. Watson, F. Mazzola, O. J. Clark, K. Underwood, I. Marković, M. McLaren, A. Hunter, D. M. Burn, L. B. Duffy, S. Barua, G. Balakrishnan, F. Bertran, P. Le Fèvre, T. K. Kim, G. van der Laan, T. Hesjedal, P. Wahl, and P. D. C. King, Nano Letters 18, 4493 (2018).
@@
- Duvjir et al. (2018) G. Duvjir, B. K. Choi, I. Jang, S. Ulstrup, S. Kang, T. Thi Ly, S. Kim, Y. H. Choi, C. Jozwiak, A. Bostwick, E. Rotenberg, J.-G. Park, R. Sankar, K.-S. Kim, J. Kim, and Y. J. Chang, Nano Letters 18, 5432 (2018).
@@
- Diego et al. (2021) J. Diego, A. H. Said, S. K. Mahatha, R. Bianco, L. Monacelli, M. Calandra, F. Mauri, K. Rossnagel, I. Errea, and S. Blanco-Canosa, Nature Communications 12, 598 (2021).
- Terashima et al. (2003) K. Terashima, T. Sato, H. Komatsu, T. Takahashi, N. Maeda, and K. Hayashi, Phys. Rev. B 68, 155108 (2003).
@@
- Strocov et al. (2012) V. N. Strocov, M. Shi, M. Kobayashi, C. Monney, X. Wang, J. Krempasky, T. Schmitt, L. Patthey, H. Berger, and P. Blaha, Phys. Rev. Lett. 109, 086401 (2012).
@@
- Fanfarillo et al. (2021) L. Fanfarillo, D. Kopić, A. Sterzi, G. Manzoni, A. Crepaldi, D. T. Payne, W. Bronsch, V. Tsurkan, D. Croitori, J. Deisenhofer, F. Parmigiani, M. Capone, and F. Cilento, Journal of Electron Spectroscopy and Related Phenomena 250, 147090 (2021).
@@
- Peli et al. (2020) S. Peli, D. Puntel, D. Kopic, B. Sockol, F. Parmigiani, and F. Cilento, Journal of Electron Spectroscopy and Related Phenomena 243, 146978 (2020).
- Kresse and Hafner (1993) G. Kresse and J. Hafner, Phys. Rev. B 48, 13115 (1993).
- Kresse and Furthmüller (1996a) G. Kresse and J. Furthmüller, Computational Materials Science 6, 15 (1996a).
- Kresse and Furthmüller (1996b) G. Kresse and J. Furthmüller, Phys. Rev. B 54, 11169 (1996b).
- Blöchl (1994) P. E. Blöchl, Phys. Rev. B 50, 17953 (1994).
- Kresse and Joubert (1999) G. Kresse and D. Joubert, Phys. Rev. B 59, 1758 (1999).
- Perdew et al. (1996) J. P. Perdew, K. Burke, and M. Ernzerhof, Phys. Rev. Lett. 77, 3865 (1996).
@@
- Jolie et al. (2019) W. Jolie, T. Knispel, N. Ehlen, K. Nikonov, C. Busse, A. Grüneis, and T. Michely, Phys. Rev. B 99, 115417 (2019).
- Pizzi et al. (2020) G. Pizzi, V. Vitale, R. Arita, S. Blügel, F. Freimuth, G. Géranton, M. Gibertini, D. Gresch, C. Johnson, T. Koretsune, J. Ibañez-Azpiroz, H. Lee, J.-M. Lihm, D. Marchand, A. Marrazzo, Y. Mokrousov, J. I. Mustafa, Y. Nohara, Y. Nomura, L. Paulatto, S. Poncé, T. Ponweiser, J. Qiao, F. Thöle, S. S. Tsirkin, M. Wierzbowska, N. Marzari, D. Vanderbilt, I. Souza, A. A. Mostofi, and J. R. Yates, Journal of Physics: Condensed Matter 32, 165902 (2020).
@@
- Falke et al. (2021) Y. Falke, N. Ehlen, G. Marini, A. V. Fedorov, V. Y. Voroshnin, B. V. Senkovskiy, K. Nikonov, M. Hoesch, T. K. Kim, L. Petaccia, G. Di Santo, T. Szkopek, G. Profeta, and A. Grüneis, Phys. Rev. B 104, 235137 (2021).
- Sugimoto et al. (2015) T. Sugimoto, D. Ootsuki, C. Morice, E. Artacho, S. S. Saxena, E. F. Schwier, M. Zheng, Y. Kojima, H. Iwasawa, K. Shimada, M. Arita, H. Namatame, M. Taniguchi, M. Takahashi, N. L. Saini, T. Asano, R. Higashinaka, T. D. Matsuda, Y. Aoki, and T. Mizokawa, Phys. Rev. B 92, 041113 (2015).
- Moser (2017) S. Moser, Journal of Electron Spectroscopy and Related Phenomena 214, 29 (2017).
- Henke et al. (2020) J. Henke, F. Flicker, J. Laverock, and J. van Wezel, SciPost Phys. 9, 56 (2020).
@@
- Tamura et al. (1985) E. Tamura, R. Feder, J. Krewer, R. Kirby, E. Kisker, E. Garwin, and F. King, Solid State Communications 55, 543 (1985).
- Claessen et al. (1990) R. Claessen, I. Schafer, and M. Skibowski, Journal of Physics: Condensed Matter 2, 10045 (1990).
- Yilmaz and Vescovo (2022) T. Yilmaz and E. Vescovo, Phys. Rev. B 105, L041101 (2022).
- Wang et al. (2015) H. Wang, C. Zhang, and F. Rana, Nano Letters 15, 339 (2015).
@@
- Cilento et al. (2018) F. Cilento, G. Manzoni, A. Sterzi, S. Peli, A. Ronchi, A. Crepaldi, F. Boschini, C. Cacho, R. Chapman, E. Springate, H. Eisaki, M. Greven, M. Berciu, A. F. Kemper, A. Damascelli, M. Capone, C. Giannetti, and F. Parmigiani, Science Advances 4, eaar1998 (2018).